OpenStudy (jenniferjuice):

Solve x2 - 7x = -13.

OpenStudy (anonymous):

x squared?

OpenStudy (anonymous):

So first, rewrite it as a quadratic

OpenStudy (anonymous):

So: x^2 - 7x + 13 = 0

OpenStudy (anonymous):

then try to factor or use the quadratic formula

OpenStudy (jenniferjuice):

how do i know what is a b and c?

OpenStudy (jenniferjuice):

@Nairz

hero (hero):

\(\color\red{a}x^2 + \color\green{b}x + \color\orange{c} = 0\) \(\color\red{1}x^2 + \color\green{-7}x + \color\orange{13} = 0\)

OpenStudy (jenniferjuice):

i think its either a or b

hero (hero):

Please show your work

hero (hero):

Did you use this formula: \(x = \dfrac{-b \pm \sqrt{b^2 - 4ac}}{2a}\)

OpenStudy (jenniferjuice):